A random variable X has the following probability mass function X 2 3 4 5 6 7 8 P(X=x) 1/16 2k 3/16 1/4 3/16 1/8 k i) Find k, hence determine the mean of the distribution. ii) Obtain �3≤�5) iii) Determine F(5)

1
Expert's answer
2021-11-29T13:09:00-0500

i)


"\dfrac{1}{16}+2k+\dfrac{3}{16}+\dfrac{1}{4}+\dfrac{3}{16}+\dfrac{1}{8}+k=1"

"3k=\dfrac{3}{16}"

"k=\dfrac{1}{16}"

"E(X)=2(\dfrac{1}{16})+ 3(\dfrac{2}{16})+ 4(\dfrac{3}{16})+ 5(\dfrac{4}{16})"

"+ 6(\dfrac{3}{16})+ 7(\dfrac{2}{16})+ 8(\dfrac{1}{16})=5"

ii)

"P(3\leq X\leq 5)=P(X=3)+P(X=4)"

"+P(X=5)=\dfrac{2}{16}+ \dfrac{3}{16}+ \dfrac{4}{16}=\dfrac{9}{16}"

iii)


"F(5)=f(2)+f(3)+f(4)+f(5)"

"=\dfrac{1}{16}+\dfrac{2}{16}+ \dfrac{3}{16}+ \dfrac{4}{16}=\dfrac{5}{8}"